# Forgetting is Everywhere

### v1

@@ 5.2 Approximate learners can benefit from forgetting
Takeaway 5.1 Forgetting is ubiquitous in deep learning. The trade-off between training efficiency and forgetting determines the optimal amount to forget–in deep learning, this is rarely zero.
@@ 5.3 Implications of distribution shift
Distribution shift and stochasticity strongly influence forgetting dynamics. In i.i.d. environments, the interaction process is stationary and induced futures evolve smoothly. When learner hyperparameters are well-tuned and training conditions are stable, forgetting dynamics become smooth and predictable, allowing the learner to balance adaptation and retention effectively. In such settings, training efficiency is high because the learner’s updates operate under consistent conditions, and the underlying data distribution provides consistent feedback for improvement.
Figure 4: Approximate learners can benefit from non-zero forgetting. We analyse how training efficiency and forgetfulness co-vary across learning algorithms. Training efficiency is measured as the inverse of the normalised area under the training loss curve, an approximate but informative proxy for learning speed and convergence quality in the setting we study. Forgetfulness is quantified as the mean 40-step propensity to forget, $\bar{\Gamma}_{40}(t)$, across training steps. Left: In a regression task, varying stochastic gradient descent’s momentum parameter shows that higher momentum increases forgetfulness, with maximum training efficiency at $0.9$ momentum. Right: Varying model size shows that maximum efficiency occurs at 20 parameters. In both cases, the most efficient learners exhibit some forgetting–too little slows adaptation, too much destabilises learning–highlighting a fundamental trade-off.
In CL, abrupt shifts in the observation distribution cause discontinuous changes in $Z_{t}$ and its induced futures. Consequently, the magnitude of consistency violation abruptly increases at task boundaries (Figure 3) as the learner must rapidly adapt to a new task.
RL, however, remains non-stationary throughout the training process. The learner’s data distribution evolves in response to its policy, creating feedback loops that continually perturb its state. This interplay between environmental and algorithmic non-stationarity induces chaotic fluctuations in the forgetting profile (Figure 5). Current neural network optimisation algorithms, designed for the more stationary supervised learning setting, struggle to stabilise these dynamics. As a result, RL learners often fail to maintain a consistent balance between retention and adaptation, leading to poor learning dynamics and reduced training efficiency.
Takeaway 5.2 Stable forgetting dynamics improve learning efficiency by promoting a balanced trade-off between adaptation and retention.
Figure 5: Chaotic forgetting dynamics in RL. We show TD loss and Q-value evaluation across ten seeds for a DQN learner trained on the cartpole environment. While neural networks trained on supervised learning can exhibit fluctuating forgetting dynamics, when hyperparameters are well-tuned, their dynamics are typically smooth and self-stabilising. In contrast, neural networks trained via RL display persistent, chaotic fluctuations driven by non-stationarity in the environment. Current RL algorithms poorly regulate these dynamics, resulting in unstable learning and reduced efficiency.
@@ 6 Conclusion
In this work, we have provided a conceptual foundation for understanding forgetting in general learning systems from a predictive perspective. We showed that forgetting can be understood as a violation of self-consistency in a learner’s predictive distribution and formalised the conditions under which forgetting occurs. We proposed an operational measure, the learner’s propensity to forget (Definition 4.2), which allowed us to validate this formalism in practice. Our results demonstrate that measures derived from the predictive formulation faithfully capture the extent of forgetting in practice and provide a clear distinction between constructive and destructive adaptation.
Empirically, we found that forgetting is pervasive across diverse settings and is determined by the interaction between the learner and the environment. Furthermore, we identified a trade-off between forgetting and training efficiency, with CL and RL settings particularly susceptible to forgetting and thus exhibiting poorer training efficiency than other learning settings.
By providing a conceptual formalism of forgetting, our work offers a principled foundation for analysing and understanding forgetting. We hope this perspective will guide the development of learning algorithms that can adapt effectively without sacrificing prior abilities, and provide a better conceptual basis for reasoning about how a learner’s capabilities emerge, persist, and degrade over time.
@@ Acknowledgements
The authors would like to thank Guy Frankel, Jonathon Hare, Henry Gouk, Andras Szecsenyi and Matej Sandor for their help and valuable insights. This research was enabled by computer resources provided by the School of Informatics at the University of Edinburgh. Funding for this research was provided by EPSRC through a PhD studentship within the CDT in Machine Learning Systems hosted in the School of Informatics, University of Edinburgh (EP/Y03516X/1).
@@
The appendix is organised as follows: Appendix A defines the notation, Appendix B expands on the theory, Appendix C gives illustrative thought experiments, Appendix D provides additional details regarding empirical implementation, Appendix E reports additional results, and Appendix F details our experiments.
@@ Appendix A Notation
$\mathcal{H}$ Set of histories $\bigcup_{t=0}^{\infty}(\mathcal{X}\times\mathcal{Y})^{t}$
@@
$H_{0:t}$ History up to time $t$ Sequence $((X_{1},Y_{1}),\dots,(X_{t},Y_{t}))$
@@
$(e,p_{X_{1}})$ Environment $p_{X_{1}}\in\mathcal{D}(\mathcal{X})$,  $e:\mathcal{H}\to\mathcal{D}(\mathcal{X})$
@@
$q_{e}(\cdot\mid h,y)$ Learner-sampled hybrid input distribution given history $h$ and output $y$ $\mathcal{D}(\mathcal{X})$
@@ A.1 Framework generality
Table 2: This table provides specific examples of values for the abstract variables: $X_{t},Y_{t}$ and $Z_{t}$ across different common machine learning settings.
@@ Appendix B Theoretical results
B.1 One-step consistency condition
After the update, a non-forgetting learner satisfies the following conditional independence rules:
@@ B.1 One-step consistency condition
Furthermore, the environment admits the following independence:
[…]
The one-step consistency operator is obtained by marginalising over observations $X_{t},Y_{t}$ and updated state $Z_{t}$:
$q^{*}(H^{t+1:\infty}\mid Z_{t-1},H_{0:t-1})=\int q(H^{t+1:\infty},X_{t},Y_{t},Z_{t}\mid Z_{t-1},H_{0:t-1})\,\mathrm{d}X_{t}\,\mathrm{d}Y_{t}\,\mathrm{d}Z_{t}.$ (15)
This can factorised, with the conditional independence rules in (13) and (14) applied, to derive the following:
$=\int q(H^{t+1:\infty}\mid Z_{t},H_{0:t})\,u(Z_{t}\mid Z_{t-1},X_{t},Y_{t})$
$\qquad\qquad q_{e}(X_{t}\mid H_{0:t-1},Y_{t})\,q_{f}(Y_{t}\mid Z_{t-1},X_{t-1})\,\mathrm{d}X_{t}\,\mathrm{d}Y_{t}\,\mathrm{d}Z_{t}.$
which is the 1-step consistency condition from (7):
$q(H^{t+1:\infty}\mid Z_{t-1},H_{0:t-1})=\mathbb{E}_{Y_{t},X_{t},Z_{t}}\left[q(H^{t+1:\infty}\mid Z_{t},H_{0:t})\right],$ (17)
@@
$Y_{t}\sim q_{f}(\cdot\mid Z_{t-1},X_{t-1}),\quad X_{t}\sim q_{e}(\cdot\mid H_{0:t-1},Y_{t}),\quad Z_{t}\sim u(\cdot\mid Z_{t-1},X_{t},Y_{t}).$
B.2 k-step consistency condition
We can apply the same reasoning to derive the $k$-step consistency condition. Denoting $t^{\prime}$ as $t^{\prime}=t+k-1$, we get:
$q_{k}^{*}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})$ (18)
@@
$\qquad q_{e}(X_{s}\mid H_{0:s-1},Y_{s})q_{f}(Y_{s}\mid Z_{s-1},X_{s-1})\big]\mathrm{d}X_{t:t^{\prime}}\,\mathrm{d}Y_{t:t^{\prime}}\,\mathrm{d}Z_{t:t^{\prime}},$
which is the $k$-step consistency condition from Definition 4.1:
$q_{k}^{*}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})=\mathbb{E}_{Y_{t:t^{\prime}},X_{t:t^{\prime}},Z_{t:t^{\prime}}}\left[q(H^{t+k:\infty}\mid Z_{t^{\prime}},H_{t^{\prime}})\right].$ (19)

### v2

@@ 5.3 Approximate learners can benefit from forgetting
Takeaway 3.
Forgetting is ubiquitous in deep learning. The trade-off between training efficiency and forgetting determines the optimal amount to forget–in deep learning, this is rarely zero.
Figure 4: Approximate learners can benefit from non-zero forgetting. We analyse how training efficiency and forgetfulness co-vary across learning algorithms. Training efficiency is measured as the inverse of the normalised area under the training loss curve, an approximate but informative proxy for learning speed and convergence quality in the setting we study. Forgetfulness is quantified as the mean 40-step propensity to forget, $\Gamma_{40}(t)$, across training steps. Left: In a regression task, varying stochastic gradient descent’s momentum parameter shows that higher momentum increases forgetfulness, with maximum training efficiency at $0.9$ momentum. Right: Varying model size shows that maximum efficiency occurs at 20 parameters. In both cases, the most efficient learners exhibit some forgetting–too little slows adaptation, too much destabilises learning–highlighting a fundamental trade-off.
@@ 5.4 Implications of distribution shift
Figure 5: DQN actively manages the information acquisition-retention trade-off. We show TD loss, Q-value evaluation, and the forgetting profile of a DQN learner trained on cartpole across ten seeds. Early in training, TD loss is low; it rises as the agent acquires new information, then decreases once knowledge is integrated. The forgetting curve follows this trajectory, highlighting that forgetting is a deliberate mechanism for balancing knowledge acquisition with knowledge retention.
Distribution shift and stochasticity strongly influence forgetting dynamics. In i.i.d. environments, the interaction process is stationary, and the predictive distributions are stable. When learner hyperparameters are well-tuned and training conditions are stable, the learner effectively balances adaptation and retention. In such settings, training efficiency is high because the learner’s updates operate under consistent conditions, and the underlying data distribution provides consistent feedback for improvement.
In CL, abrupt shifts in the observation distribution cause discontinuous changes in state, $Z_{t}$, and the predictive distribution. Consequently, the magnitude of self-consistency violation abruptly increases at task boundaries (Figure 3) as the learner must rapidly adapt to a new task.
All learning involves balancing the integration of new information with the retention of current knowledge. RL presents this challenge in an extreme form. Here, the learner’s policy influences future observations, inducing continual non-stationarity. In DQN, for example, as the agent experiences new transitions, the TD loss rises because the agent incorporates new information (Figure 5). As the agent consolidates this information, the TD loss declines and the rate of information acquisition plateaus. The forgetting curve follows the TD loss because forgetting information is the mechanism by which the agent manages this process, demonstrating that forgetting is an essential component of RL.
Takeaway 4.
Forgetting is an integral component of learning: effective learning requires selectively forgetting outdated knowledge to integrate new information.
@@ 6 Conclusion
In this work, we proposed a general, algorithm- and task-agnostic formulation of forgetting, characterising it as the temporal inconsistency of a learner’s predictive distribution. This perspective provides a unified conceptualisation of forgetting, disentangling it from backward transfer and separating it from parameter updates. Furthermore, we introduce the propensity to forget as an operational measure of forgetting (§ 4.2), and under Assumptions B.3, B.2, B.4, B.1 and B.5, we prove that Bayesian learners do not forget. This result confirms the intuition that some learners can adapt to new data without forgetting.
Our empirical analysis across a diverse set of algorithms and task settings shows that forgetting is widespread in deep learning and is influenced by the interactions between the learner and the environment. Furthermore, we observe that training efficiency and forgetting are not monotonic: in the settings we study, an intermediate amount of forgetting can maximise learning efficiency. This demonstrates the importance of considering forgetting when designing and evaluating learning algorithms.
Looking ahead, several directions for future work remain. These include identifying non-forgetful learners that are not Bayesian and exploring how increasing the self-consistency of learners affects learning. This would provide valuable insights into the dynamics of adaptive learning systems.
Overall, our work reframes forgetting as a fundamental property of learning, rather than a failure mode restricted to continual learning or non-stationary settings. We hope our work establishes a principled foundation for analysing how learning algorithms acquire, maintain, and lose capabilities over time, guiding the development of new learning algorithms that can optimally adapt their beliefs.
@@ Acknowledgements
The authors would like to thank Guy Frankel, Jonathon Hare, Henry Gouk, Joseph Mellor, Andras Szecsenyi, and Matej Sandor for their help and valuable insights. This research was enabled by computer resources provided by the School of Informatics at the University of Edinburgh. Funding for this research was provided by EPSRC through a PhD studentship within the CDT in Machine Learning Systems hosted in the School of Informatics, University of Edinburgh (EP/Y03516X/1).
@@
The appendix is organised as follows: § A defines the notation, § B provides a formal theoretical exposition, § C gives illustrative thought experiments, § D provides additional details regarding empirical implementation, § E reports additional results, and § F details our experiments.
@@ Appendix A Notation
$\mathcal{H}$ Set of histories $\bigcup_{t=0}^{\infty}(\mathcal{X}\times\mathcal{Y})^{t+1}$
@@
$H_{0:t}$ History up to time $t$ Sequence $((X_{0},Y_{0}),\dots,(X_{t},Y_{t}))$
@@
$(e,p_{X_{0}})$ Environment $e:\mathcal{H}\to\mathcal{D}(\mathcal{X}),\;p_{X_{0}}\in\mathcal{D}(\mathcal{X})$
@@
$q_{e}(\cdot\mid h,y)$ Learner-sampled input distribution given history $h$ and output $y$ $\mathcal{D}(\mathcal{X})$
@@ A.1 Framework generality
Table 2: This table provides specific examples of values for the abstract variables: $X_{t},Y_{t}$, and $Z_{t}$ across different common machine learning settings.
@@ Appendix B Theoretical results
B.1 One-step simulated marginalisation
The one-step simulated marginalisation is obtained by marginalising over observations $X_{t},Y_{t}$, and the updated state $Z_{t}$:
$q(H^{t+1:\infty}\mid Z_{t-1},H_{0:t-1})=\int q(H^{t+1:\infty},X_{t},Y_{t},Z_{t}\mid Z_{t-1},H_{0:t-1})\,\mathrm{d}X_{t}\,\mathrm{d}Y_{t}\,\mathrm{d}Z_{t}.$ (13)
This can be factorised to:
$=\int q(H^{t+1:\infty}\mid X_{t},Y_{t},Z_{t},Z_{t-1},H_{0:t-1})\,u(Z_{t}\mid Z_{t-1},X_{t},Y_{t},H_{0:t-1})$
$\qquad\qquad\cdot q_{e}(X_{t}\mid Y_{t},Z_{t-1},H_{0:t-1})\,q_{f}(Y_{t}\mid Z_{t-1},H_{0:t-1})\,\mathrm{d}X_{t}\,\mathrm{d}Y_{t}\,\mathrm{d}Z_{t}.$
In generic learning scenarios and under the assumption of learning over time, we assume no access to the history. Thus, for non-forgetting algorithms, the state update, input, and prediction are conditionally independent of the history. Furthermore, the induced futures distribution is conditionally independent of past learner states:
@@ B.1 One-step simulated marginalisation
This reduces the factorised form to:
$=\int q(H^{t+1:\infty}\mid Z_{t},H_{0:t})\,u(Z_{t}\mid Z_{t-1},X_{t},Y_{t})$ (16)
$\qquad\qquad\cdot q_{e}(X_{t}\mid Y_{t},Z_{t-1})\,q_{f}(Y_{t}\mid Z_{t-1},X_{t-1})\,\mathrm{d}X_{t}\,\mathrm{d}Y_{t}\,\mathrm{d}Z_{t}.$
which is the 1-step simulated marginalisation:
$q^{*}_{1}(H^{t+1:\infty}\mid Z_{t-1},H_{0:t-1})=\mathbb{E}_{X_{t},Y_{t},Z_{t}}\left[q(H^{t+1:\infty}\mid Z_{t},H_{0:t})\right],$ (17)
@@
$X_{t}\sim q_{e}(\cdot\mid Y_{t},Z_{t-1}),\quad Y_{t}\sim q_{f}(\cdot\mid Z_{t-1},X_{t-1}),\quad Z_{t}\sim u(\cdot\mid Z_{t-1},X_{t},Y_{t}).$
B.2 k-step simulated marginalisation
We can apply the same reasoning to derive the $k$-step simulated marginalisation. If we denote $t^{\prime}$ as $t^{\prime}=t+k-1$, we get:
$q(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})$ (18)
@@
$\qquad\cdot q_{e}(X_{s}\mid Y_{s},Z_{s-1})q_{f}(Y_{s}\mid Z_{s-1},X_{s-1})\big]\mathrm{d}X_{t:t^{\prime}}\,\mathrm{d}Y_{t:t^{\prime}}\,\mathrm{d}Z_{t:t^{\prime}},$
which is the $k$-step simulated marginalisation given in § 4.2:
$q_{k}^{*}(H^{t+k:\infty}\mid Z_{t-1},H_{0:t-1})=\mathbb{E}_{X_{t:t^{\prime}},Y_{t:t^{\prime}},Z_{t:t^{\prime}}}\left[q(H^{t+k:\infty}\mid Z_{t^{\prime}},H_{0:t^{\prime}})\right].$ (19)